# Mapping the Spiral of Silence: Surveying Unspoken Opinions in Online Communities

### v1

@@ Abstract
We often treat social media as a lens onto society. How might that lens be distorting the actual popularity of political and social viewpoints? In this paper, we examine the difference between the viewpoints publicly posted in a community and the privately surveyed viewpoints of community members, contributing a measurement of a theory called the “spiral of silence.” This theory observes that people are less likely to voice their opinion when they believe they are in the minority—leading to a spiral where minority opinions are less likely to be shared, so they appear even further in the minority, and become even less likely to be shared. We surveyed active members of politically oriented Reddit communities to gauge their willingness to post on contentious topics, yielding 627 responses from 108 participants about 11 topics and 33 subreddits. We find that 72.6% of participants who perceive themselves in the minority remain silent, and are only half as likely to post their viewpoint compared to those who believe their opinion is in the majority. Communities perceived as being more inclusive reduce the magnitude of this effect. These results emphasize how far out of step the opinions we see online may be with the population they purport to represent.
1 Introduction
@@
Content shared on social media is often treated as a reflection of broader public opinion (Anstead and O’Loughlin 2015; McGregor 2020; Asur and Huberman 2010). But is this view distorted? For example, in 2019, a racist photo of the governor of Virginia surfaced online. As reported in The New York Times, this event led to universal condemnation of the governor on social media, leading major politicians to demand that he resign (Cohn and Quealy 2019). However, a poll of Virginians showed that most people, and especially Black constituents, actually wanted the governor to remain in office (Jamison and Clement 2019). Such voices were less vocal and less visible on social media, despite their prevalence.
@@
Divergence between what we see on social media and what the public believes is not uncommon. What causes this distortion? While this could be an artifact of users’ preferences for posting online on certain topics, or perhaps a product of social loafing (Karau and Williams 1993), we explore an alternative: users may feel uncomfortable sharing their opinion if they feel they are in the minority. We ground this explanation in the spiral of silence theory (Noelle-Neumann 1974). The theory posits that, due to a fear of social isolation, people may self-silence, defined as deciding not to share their opinion, when they view themselves to be in the minority. Self-silencing leads to less of the opinion being expressed in the group, making the viewpoint appear even further minoritized, producing a downward spiral where people self-silence even further. Since being proposed, the spiral of silence has been empirically studied in face-to-face communication across multiple issue types and participant populations (Scheufle and Moy 2000; Matthes, Knoll, and von Sikorski 2018). As online communication forms like social media continue to proliferate, it raises the question how the spiral of silence might affect user behavior in this setting.
@@
Measuring the spiral of silence on social media, especially across diverse online communities, is challenging. Not all topics are likely to trigger the spiral of silence (Lee et al. 2004), and identifying such topics may require insider knowledge of community norms and practices (Chen and Berger 2013; Hessel and Lee 2019). Furthermore, even with a set of topics in hand, the process of measuring rates of self-silencing requires capturing user attitudes, beliefs, and behaviors that cannot be observed from online content itself. This traditional approach of analyzing existing posts online is self-defeating, since it excludes viewpoints that users feel uncomfortable sharing online.
@@
In our work, we develop a human-plus-algorithm pipeline to generate topics that are contextually relevant to a community but also likely to spark internal disagreement. Then, drawing on survey methodologies common in empirical studies of the spiral of silence (Matthes, Knoll, and von Sikorski 2018), we measure whether the spiral of silence is occurring on those topics in those communities. Further, we evaluate whether online community factors such as moderation and inclusivity mitigate or amplify the spiral. We apply this method to study the spiral of silence on Reddit — a popular social media platform in the United States — across a number of subreddits focused on political and social issues but with varying norms and values.
@@
We find that those who believe their opinion is in the minority remain silent $72.6\%$ of the time, and these opinions are only half as likely to be posted compared to those in the majority. They are also less likely to upvote posts sharing their opinion. However, this self-silencing effect is mitigated by the community’s perceived level of inclusivity. We test these results through a series of three mixed-effects models and find that higher inclusivity is not only positively associated with posting overall, but leads to $2.3$ times more of an increase in opinion expression for those who believe they are in the minority.
@@
Our findings reinforce how social media data can present a substantial distortion of broader public opinion, systematically under-representing minority viewpoints. There are certainly some topics where minority opinions are normatively undesirable to represent, even if a portion of the population agrees with them, such as, hate speech or misinformation. However, for many other topics where the minority voice instead reflects under-represented groups or moderate views from people who believe themselves to be in the minority, communities may want to mitigate the spiral and hear from a more representative set of voices (Farjam and Loxbo 2024; Bowen and Blackmon 2003). To reduce rates of self-silencing, we discuss tangible actions that our results suggest online moderators and platform designers can take to increase perceived community inclusivity and mitigate the spiral of silence.
2 Background
@@
Social media is often viewed as a mirror of society, or at the very least, one of our best available instruments for understanding public opinion. Social media data has been used for a range of tasks from forecasting box-office performances (Asur and Huberman 2010) to understanding political support for candidates (McGregor 2020). Furthermore, social media content can actively shape people’s opinions and influence their actions (Laranjo et al. 2015).
@@
However, what we see online may not be a complete or accurate picture of public opinion. This is partly because online content originates from a small fraction of platform users (Ruths and Pfeffer 2014). The majority of online communities consist of “lurkers” or individuals who visit the community but choose to remain silent (Sun, Rau, and Ma 2014). There are many reasons why people may choose to stay as lurkers. In some cases, users may just want to learn about a topic or are not motivated to post if they see their opinion reflected in existing content (Nonnecke and Preece 2001). Other times, however, users may want to share their opinions but are uncomfortable doing so. The disproportionate silencing of certain viewpoints skews the content that we see online away from the true opinions of users.
@@
The spiral of silence is an empirical effect, not an unavoidable constant applying to all people: there are “avant-garde” and “hard core” individuals who stand up for their viewpoints even when they are in the minority (Noelle-Neumann 1977). Through empirical studies, work has proposed factors that modulate the spiral of silence. Individual-level antecedents, including personality traits (Hayes, Glynn, and Shanahan 2005; Neuwirth, Frederick, and Mayo 2007), opinion certainty (Matthes, Knoll, and von Sikorski 2018), and issue type (Gearhart and Zhang 2018) can influence opinion expression. In social media settings, prior work has focused on identifying platform affordances that influence self-silencing behavior (Wu and Atkin 2018; Pang et al. 2016). For example, greater perceived anonymity can dampen the effects as the potential social consequences of speaking out are diluted (Wu and Atkin 2018). Even commenting versus liking content can impact the degree of self-silencing (Pang et al. 2016).
While we have insight into how platform-level design decisions influence self-silencing, platforms can house many different online communities each with their own rules, norms, and values. Given the unique configuration of each online community, it is not surprising that users’ behaviors, such as conversation patterns (Choi et al. 2015) or self-disclosure (Yang, Yao, and Kraut 2017), will differ across communities. At the moment, the impact of these community-level design decisions has not been factored into our understanding of how the spiral of silence manifests. This leads us to our research questions: {addmargin}[1.25em]0em RQ1. How does users’ opinion expression on Reddit differ when they perceive their viewpoint as being in the majority versus the minority?
RQ2. How do online community design factors influence opinion expression of incongruent viewpoints?
3 Factors Influencing Opinion Expression
@@ 3.1 Opinion Incongruency
The spiral of silence theory provides a potential explanation for opinion expression on social media platforms. Previous studies (Hampton et al. 2014; Kushin, Yamamoto, and Dalisay 2019) on social media platforms found evidence that users are less likely to express incongruent viewpoints. We expect a similar phenomenon on Reddit. Since subreddits offer a sense of community and socializing opportunities (Moore and Chuang 2017), community members may still experience a fear of social isolation. Repercussions, such as being downvoted or facing personal attacks, can trigger this fear (Neubaum and Krämer 2018; Cheng et al. 2017). These consequences encourage users to self-silence when they have an opinion they believe is in the minority. Thus, we expect our study will yield results in accordance with the spiral of silence: {addmargin}[1.25em]0em Hypothesis 1: Likelihood of opinion expression is positively associated with viewpoint congruency.
3.2 Community Inclusion and Diversity
How online communities are designed will influence how users behave, especially regarding sharing their opinions publicly. A greater sense of belonging can lead users to be more active as they feel more included and trusting of others in the group. Similarly, feeling psychologically safe can foster confidence and empower users to voice their opinions (Edmondson and Lei 2014).
Feeling a sense of belonging is an essential human need that applies not only to offline but also online settings (Maslow 1958; Lambert et al. 2013). Hagerty et al. (1992) proposed two key dimensions that define a sense of belonging: (1) having involvement in a group be valued, and (2) feeling a sense of fit or congruency with other group members. When these criteria are met, people are more likely to be involved in a group and find this involvement to be more meaningful (Hagerty et al. 1996). The same principles apply to belonging in online communities (Lampe et al. 2010). Users with strong shared identities or bonds are more likely to be active in online groups (Ren, Kraut, and Kiesler 2007). When people feel included, they are more likely to have a sense of familiarity and trust other community members (Zhao et al. 2012). This encourages users to share more freely as they feel less at risk of being judged when sharing their viewpoints. Thus, we expect that:
{addmargin}
[1.25em]0em Hypothesis 2a: The perceived inclusivity of a community is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
Who is part of the community also matters. Users frequently mention the fear of being personally attacked as a primary reason for not posting minority opinions (Neubaum and Krämer 2018). In both online and offline settings, existing research has found that people are more secure when sharing ideas in psychologically safe communities (Newman, Donohue, and Eva 2017). That is, when people foresee fewer negative interpersonal consequences, they are more willing to take potentially controversial actions. Community diversity is an important antecedent for psychological safety (Newman, Donohue, and Eva 2017). Emphasizing diversity can foster a positive climate in which community members, regardless of their background, more strongly identify as members of the group and feel psychologically safe (McKay et al. 2007). In this setting, group members are more willing to engage in potentially risky behaviors. From this literature, we hypothesize the following:
{addmargin}
@@ 3.2 Community Inclusion and Diversity
[1.25em]0em Hypothesis 2b: The diversity of community members is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
To conceptualize inclusion and diversity, we draw on work that has sought to quantify the values of online communities (Weld, Zhang, and Althoff 2022). We utilize the instrument proposed in this work (Weld, Zhang, and Althoff 2024) to capture users’ perceived sense of inclusion and diversity across different subreddits.
3.3 Community Moderation
Moderation practices frequently shape the dynamics of online communities. Moderators can set the tone of the online community by encouraging other users to imitate their actions and by curbing instances of behavior that they deem to fall outside community norms (Seering, Kraut, and Dabbish 2017). Removing or prescreening content that does not adhere to the community’s norms is an effective way of reducing “bad behavior” and can positively impact community health (Kraut and Resnick 2012; Seering, Kraut, and Dabbish 2017). By filtering out potentially harassing or offensive submissions, moderators can assuage members’ fear of being personally attacked and facilitate more open opinion-sharing (Gibson 2019). For example, Wadden et al. (2021) studied mental health discussions in moderated and unmoderated spaces. They found that users were not only more likely to post in moderated spaces but also that they were more willing to share vulnerable messages in these discussions. From these results, we expect that users feel more comfortable sharing minoritized opinions in communities with more active moderation. Thus, we hypothesize:
{addmargin}
@@ 3.3 Community Moderation
[1.25em]0em Hypothesis 3: Moderation activity is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.

### v2

@@ Abstract.
We often treat social media as a lens onto society. How might that lens distort the popularity of political and social viewpoints? We examine discrepancies between publicly posted and privately surveyed opinions within communities, contributing a measurement of the “spiral of silence” theory; the theory posits people are less likely to voice opinions when they believe they hold minority views, creating a reinforcing cycle where these opinions are expressed less. We surveyed members of politically-oriented Reddit communities about their willingness to post on contentious topics, yielding 439 responses across twelve subreddits. 72.1% of participants who perceive themselves in the minority remain silent and are half as likely to post compared to those who believe their opinion is in the majority. Community design factors, such as perceived diversity, are associated with less self-silencing. We provide recommendations for counteracting self-silencing at the community level (e.g., positive reinforcement, more transparent moderation). Overall, these results reveal gaps between online discourse and broader public opinion.
spiral of silence, online communities
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790361isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computing
Figure 1. The political orientation of content posted on r/politics is not representative of its members’ ideologies. Since submissions on r/politics must contain URLs to articles from approved news and media organizations (pol, [n. d.]), we analyze the political leanings of the linked-to news sources. Compared to the surveyed political orientations of r/politics members with the ideologies expressed in the links shared to the subreddit, there is $1.3$ times more liberal-leaning links than active liberal members on r/politics (see Appendix A.1 for methodological details).
Comparison between the political orientation of content posted on r/politics and the surveyed political ideologies of its active members from Prolific. Comparison between the political orientation of content posted on r/politics and the surveyed political ideologies of its active members from Prolific. The figure shows that 74.4% of shared links come from liberal-leaning sources, while only 58.6% of active members identify as liberal, revealing a distortion in representation.
1. Introduction
@@
Content shared on social media is often treated as a reflection of broader public opinion (Anstead and O’Loughlin, 2015; McGregor, 2020; Asur and Huberman, 2010). But is this view distorted? For example, in 2019, a racist photo of the governor of Virginia surfaced online. As reported in The New York Times, this event led to universal condemnation of the governor on social media, leading major politicians to demand that he resign (Cohn and Quealy, 2019). However, a poll of Virginians showed that most people, and especially Black constituents, were more in favor of the governor remaining in office (Jamison and Clement, 2019). Such voices were less vocal and less visible on social media, despite their prevalence.
@@
Divergence between what we see on social media and what the public believes is not uncommon. For example, comparing the content submitted to the subreddit r/politics — a massive online community for sharing and discussing US political news — to the surveyed ideologies of active members, there is a large distortion (details in Appendix A.1): $74.4\%$ of the content shared on the subreddit has a liberal view, compared to only $58.6\%$ of what active members self-report (Fig. 1). This distortion can arise in liberal, conservative, and mixed spaces. What causes this distortion? While this could be an artifact of users’ preferences for posting online on certain topics, or perhaps a product of social loafing (Karau and Williams, 1993), we explore an alternative: users may refrain from sharing their opinion when they perceive themselves as being in the minority. The spiral of silence theory (Noelle-Neumann, 1974) explains this behavior as a response to the fear of social isolation, where individuals self-silence — or decide not to share their opinion — rather than risk social disapproval. Self-silencing then leads to less of the opinion being expressed in the group, making the viewpoint appear even further minoritized. This reaction produces a downward spiral as people become more likely to self-silence. Since being proposed, the spiral of silence has been empirically studied in face-to-face communication across multiple issue types and participant populations (Scheufle and Moy, 2000; Matthes et al., 2018). As online communication forms like social media proliferate, it raises the question of how the spiral of silence might affect user behavior in this setting.
@@
Measuring the spiral of silence on social media, especially across diverse online communities, is challenging. Not all topics are likely to trigger the spiral of silence (Lee et al., 2004), and identifying such topics may require insider knowledge of community norms and practices (Chen and Berger, 2013; Hessel and Lee, 2019). Furthermore, even with a set of topics in hand, the process of measuring rates of self-silencing requires capturing user attitudes, beliefs, and behaviors that cannot be observed from online content itself. This traditional approach of analyzing existing posts online is self-defeating, since it excludes viewpoints that users feel uncomfortable sharing online.
@@
In our work, we develop a human-plus-algorithm pipeline to generate potential topics that community members feel are appropriate and in-bounds for a community to discuss, but which also likely to spark internal disagreement. Then, drawing on survey methodologies common in empirical studies of the spiral of silence (Matthes et al., 2018), we measure whether the spiral of silence occurs on those topics in those communities. Further, we evaluate whether online community factors such as moderation and diversity mitigate or amplify the spiral. We apply this method to study the spiral of silence on Reddit — a popular social media platform in the United States — across a number of subcommunities (“subreddits”) focused on political and social issues but with varying norms and values. Importantly, our analyses focus on topics that community members agree are acceptable to discuss within the subreddit but are still hesitant to engage with, and exclude topics considered inappropriate or in violation of community norms.
@@
We find that participants who believe their opinion is in the minority remain silent $72.1\%$ of the time, and these opinions are only half as likely to be posted compared to those in the majority. Participants are also less likely to upvote posts sharing minoritized opinions. However, we identify two community design factors that are linked to variations in the rate of self-silencing. We test these results through a series of three mixed-effects models, finding that participants report a higher likelihood of sharing minoritized opinions in subreddits they perceive as more diverse, but a lower likelihood of sharing in subreddits with more stringent content moderation. [^1] Code available at https://github.com/StanfordHCI/spiral-of-silence.
@@
Our findings illustrate how, within the politically-oriented subreddits we study, the distribution of viewpoints being shared online can misrepresent community members’ actual opinions, systematically marginalizing minority perspectives. There are certainly some topics where minority opinions are normatively undesirable to represent, even if a portion of the population agrees with them, such as, hate speech or misinformation. However, for many other topics where the minority voice instead reflects under-represented groups or moderate views from people who believe themselves to be in the minority, communities may want to mitigate the spiral and hear from a more representative set of voices (Farjam and Loxbo, 2024; Bowen and Blackmon, 2003). Based on our findings, we surface tangible actions that online moderators and platform designers can take, such as providing positive feedback on posts expressing diverse viewpoints and increasing moderation transparency, to help mitigate the spiral of silence.
2. How the Spiral of Silence Shapes Social Media
@@
Social media is often viewed as a mirror of society, or at the very least, one of our best available instruments for understanding public opinion. Social media data has been used for a range of tasks from forecasting box-office performances (Asur and Huberman, 2010) to understanding political support for candidates (McGregor, 2020). Furthermore, social media content can actively shape people’s opinions and influence their actions (Laranjo et al., 2015).
@@
However, what we see online may not be a complete or accurate picture of public opinion. This is partly because online content originates from a small fraction of platform users (Ruths and Pfeffer, 2014). The majority of online communities consist of “lurkers” or individuals who visit the community but choose to remain silent (Sun et al., 2014). There are many reasons why people may choose to stay as lurkers. In some cases, users may just want to learn about a topic or are not motivated to post if they see their opinion reflected in existing content (Nonnecke and Preece, 2001). Other times, however, users may want to share their opinions but are uncomfortable doing so. The disproportionate silencing of viewpoints may make our online environments appear more polarized than they actually are.
@@
The spiral of silence is an empirical effect, not an unavoidable constant applying to all people: there are “avant-garde” and “hard core” individuals who stand up for their viewpoints even when they are in the minority (Noelle-Neumann, 1977). Through empirical studies, work has proposed factors that modulate the spiral of silence. Individual-level antecedents, including personality traits (Hayes et al., 2005; Neuwirth et al., 2007), opinion certainty (Matthes et al., 2018), and issue type (Gearhart and Zhang, 2018) can influence opinion expression. In social media settings, prior work has focused on identifying platform affordances that influence self-silencing behavior (Wu and Atkin, 2018; Pang et al., 2016). For example, greater perceived anonymity can dampen the effects as the potential social consequences of speaking out are diluted (Wu and Atkin, 2018). Even the means by which opinions are shared, such as commenting versus liking content, can impact the degree of self-silencing (Pang et al., 2016).
While we have insight into how platform-level design decisions influence self-silencing, platforms can house many online communities each with their own rules, norms, and values. Given the unique configuration of each online community, it is not surprising that users’ behaviors, such as conversation patterns (Choi et al., 2015) or self-disclosure (Yang et al., 2017b), will differ across communities. At the moment, the impact of these community-level design decisions has not been factored into our understanding of how the spiral of silence manifests. To better understand how these design decisions are linked to self-silencing, this work focuses on Reddit, which is comprised of many distinct subreddits — each of which has its own governance structures, moderation practices, values, topical interests, and norms (Weld et al., 2024; Fiesler et al., 2018; Leibmann et al., 2025) — providing an ideal context for studying community-level design variation. This leads us to our research questions:
RQ1. How does users’ opinion expression on political subreddits differ when they perceive their viewpoint as being in the majority versus the minority?
RQ2. How do online community design factors on political subreddits influence opinion expression of incongruent viewpoints?
3. Factors Influencing Opinion Expression
@@ 3.1. Opinion Incongruency
The spiral of silence theory provides an explanation for viewpoint sharing patterns on social media platforms. Previous studies (Hampton et al., 2014; Kushin et al., 2019; Pang et al., 2016; Hoffmann and Lutz, 2017; Gearhart and Zhang, 2015; Liao et al., 2016) on social media platforms, such as Facebook and Twitter, found evidence that users are less likely to express incongruent viewpoints, or opinions that differ from what they perceive the majority of other members believe. We expect a similar phenomenon on Reddit. Since subreddits also offer a sense of community and socializing opportunities (Moore and Chuang, 2017), community members may still experience a fear of social isolation. This fear can be triggered on Reddit by potential social repercussions, such as being downvoted or facing personal attacks from others in the community (Neubaum and Krämer, 2018; Cheng et al., 2017). Especially when discussing stigmatized topics, users also have expressed concerns about exposure, including having sensitive content linked to their public profile or facing offline repercussions from people who know about their online account (Leavitt, 2015; Reagle, 2023). These consequences encourage users to self-silence when they have an opinion that they think differs from the majority. Given this evidence, we expect our study will yield results in accordance with the spiral of silence:
@@
Hypothesis 1: The likelihood of opinion expression is negatively associated with viewpoint incongruency (i.e., when users believe the majority of community members disagree with them).
3.2. Community Inclusion and Diversity
How online communities are designed will influence how users behave, especially regarding sharing their opinions publicly. A greater sense of belonging can lead users to be more active as they feel more included and trusting of others in the group. Similarly, feeling psychologically safe can foster confidence and empower users to voice their opinions (Edmondson and Lei, 2014).
Feeling a sense of belonging is an essential human need that applies not only to offline but also online settings (Maslow, 1958; Lambert et al., 2013). Hagerty et al. proposed two key dimensions that define a sense of belonging: (1) having involvement in a group be valued, and (2) feeling a sense of fit or congruency with other group members. When these criteria are met, people are more likely to be involved in a group and find this involvement to be more meaningful (Hagerty et al., 1996). The same principles apply to belonging in online communities (Lampe et al., 2010). Users with strong shared identities or bonds are more likely to be active in online groups (Ren et al., 2007). When people feel included, they are more likely to have a sense of familiarity and trust other community members (Zhao et al., 2012). This encourages users to share more freely, as they feel less at risk of being judged when sharing their viewpoints. Thus, we expect that:
@@ 3.2. Community Inclusion and Diversity
Hypothesis 2a: The perceived inclusivity of a community is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
Who is part of the community also matters. Users frequently mention the fear of being personally attacked as a primary reason for not posting minority opinions (Neubaum and Krämer, 2018). In both online and offline settings, existing research has found that people are more secure when sharing ideas in psychologically safe communities (Newman et al., 2017). That is, when people foresee fewer negative interpersonal consequences, they are more willing to take potentially controversial actions. Community diversity is an important antecedent for psychological safety (Newman et al., 2017). Emphasizing diversity can foster a positive climate in which community members, regardless of their background, more strongly identify as members of the group and feel psychologically safe (McKay et al., 2007). In this setting, group members are more willing to engage in potentially risky behaviors, such as sharing opinions that may deviate from what is popular. In a similar vein, prior work on the spiral of silence has suggested that those with more bridging social capital that connects them to heterogeneous networks on social media platforms, are more willing to speak up online (Sheehan, 2015). From this literature, we hypothesize the following:
Hypothesis 2b: The diversity of community members is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.
To conceptualize inclusion and diversity, we draw on work that has sought to quantify the values of online communities (Weld et al., 2022). We utilize the instrument proposed by Weld et al. (2024) to capture users’ perceived sense of inclusion and diversity across different subreddits. Perceived inclusion is defined as how “included and able to contribute new and existing member feel” while perceived diversity refers to how diverse people in the community are thought to be.
3.3. Community Moderation
Moderation practices frequently shape the dynamics of online communities. Moderators can set the tone of the online community by encouraging other users to imitate their actions and by curbing instances of behavior that they deem to fall outside community norms (Seering et al., 2017). Removing or prescreening content that does not adhere to the community’s norms is an effective way of reducing “bad behavior” and can positively impact community health (Kraut and Resnick, 2012; Seering et al., 2017). By filtering out potentially harassing or offensive submissions, moderators can assuage members’ fear of being personally attacked and facilitate more open opinion-sharing (Gibson, 2019). For example, Wadden et al. (2021) studied mental health discussions in moderated and unmoderated spaces. They found that users were not only more likely to post in moderated spaces, but also that they were more willing to share vulnerable messages in these discussions. In a setting more similar to opinion-sharing on Reddit, Wise et al. (2006) found that users were more willing to participate in a moderated online political community compared to an unmoderated space. From these results, we expect that users feel more comfortable sharing minoritized opinions in communities with more active moderation. Thus, we hypothesize:
Hypothesis 3: Moderation activity is more positively associated with the likelihood of opinion expression when users hold incongruent viewpoints.